# Cross-Field Interface-Aware Neural Operators for Multiphase Flow Simulation thanks: *Equal contribution, Corresponding author

### v1

@@ Abstract
Multiphase flow systems, with their complex dynamics, field discontinuities, and interphase interactions, pose significant computational challenges for traditional numerical solvers. While neural operators offer efficient alternatives, they often struggle to achieve high-resolution numerical accuracy in these systems. This limitation primarily stems from the inherent spatial heterogeneity and the scarcity of high-quality training data in multiphase flows. In this work, we propose the Interface Information-Aware Neural Operator (IANO), a novel framework that explicitly leverages interface information as a physical prior to enhance the prediction accuracy. The IANO architecture introduces two key components: 1) An interface-aware multiple function encoding mechanism jointly models multiple physical fields and interfaces, thus capturing the high-frequency physical features at the interface. 2) A geometry-aware positional encoding mechanism further establishes the relationship between interface information, physical variables, and spatial positions, enabling it to achieve pointwise super-resolution prediction even in the low-data regimes. Experimental results demonstrate that IANO outperforms baselines by $\sim$10% in accuracy for multiphase flow simulations while maintaining robustness under data-scarce and noise-perturbed conditions.
@@ I Introduction
Over the past few years, many deep learning techniques have been proposed to provide efficient alternatives [6, 7]. Among others, neural operators learn the infinite-dimensional mapping from functional parameters to the solution [8, 9, 10, 11], which requires training only once and predicts solutions through a single forward computation, significantly accelerating the process of solving PDEs. For instance, DeepONet [12] uses a branch network and a trunk network to process input functions and query coordinates. This architecture has been proven to approximate any nonlinear operators with a sufficiently large network. MIONet [13] extends DeepONets to solve problems with multiple input functions. Fourier neural operator (FNO) [14] operates in the spectral space using the Fast Fourier Transform, which achieves a good cost-accuracy trade-off.
@@ III-D Geometry-Aware Positional Encoding
$\mathbf{Z}_{i}=\sum_{i\in\mathcal{K}}\frac{q_{x}\cdot k_{i}}{\sum_{l\in\mathcal{K}}q_{x}\cdot k_{l}}\cdot v_{i},$ (4)
@@
$\mathbf{Z}_{i}=\sum_{i\in\mathcal{K}}\alpha_{t}(q_{x}\cdot k_{i})\cdot v_{i}=\alpha_{t}q_{x}\cdot\left(\sum_{i\in\mathcal{K}}k_{i}\otimes v_{i}\right).$ (5)
@@
$\mathbf{Z}_{i}=q_{x}+\frac{1}{M+2}\sum_{l=1}^{M+2}\alpha^{l}_{t}q_{x}\cdot\left(\sum_{i_{l}=1}^{N_{l}}k_{i_{l}}\otimes v_{i_{l}}\right),$ (6)
@@
$\mathbf{Z}^{\prime}_{i}=\text{softmax}\left(\frac{q_{x}\cdot k_{x}}{\sqrt{d_{\text{latent}}}}\right)\cdot v_{x},$ (7)
where $q_{x}$, $k_{x}$, and $v_{x}$ are computed from the updated embedding $\mathbf{Z}^{\prime}_{i}$ as follows:
$q_{x}=W_{q}\mathbf{Z}^{\prime}_{i},\quad k_{x}=W_{k}\mathbf{Z}^{\prime}_{i},\quad v_{x}=W_{v}\mathbf{Z}^{\prime}_{i}.$ (8)

### v2

@@ Abstract
Multiphase flow simulation is critical in science and engineering but incurs high computational costs due to complex field discontinuities and the need for high-resolution numerical meshes. While Neural Operators (NOs) offer an efficient alternative for solving Partial Differential Equations (PDEs), they struggle with two core challenges unique to multiphase systems: spectral bias caused by spatial heterogeneity at phase interfaces, and the persistent scarcity of expensive, high-resolution field data. This work introduces the Interface Information Aware Neural Operator (IANO), a novel architecture that mitigates these issues by leveraging readily obtainable interface data (e.g., topology and position). Interface data inherently contains the high-frequency features not only necessary to complement the physical field data, but also help with spectral bias. IANO incorporates an interface-aware function encoding mechanism to capture dynamic coupling, and a geometry-aware positional encoding method to enhance spatial fidelity for pointwise super-resolution. Empirical results across multiple multiphase flow cases demonstrate that IANO achieves significant accuracy improvements (up to $\sim$10%) over existing NO baselines. Furthermore, IANO exhibits superior generalization capabilities in low-data and noisy settings, confirming its utility for practical, data-efficient AI-based multiphase flow simulations.
@@ I Introduction
Over the past few years, many deep learning techniques have been proposed to provide efficient alternatives [6, 7, 8, 9, 10, 11]. Among others, neural operators learn the infinite-dimensional mapping from functional parameters to the solution [12, 13, 14, 15], which requires training only once and predicts solutions through a single forward computation, significantly accelerating the process of solving PDEs. For instance, DeepONet [16] uses a branch network and a trunk network to process input functions and query coordinates. This architecture has been proven to approximate any nonlinear operators with a sufficiently large network. MIONet [17] extends DeepONets to solve problems with multiple input functions. Fourier neural operator (FNO) [18] operates in the spectral space using the Fast Fourier Transform, which achieves a good cost-accuracy trade-off.
@@ III-D Geometry-Aware Positional Encoding
$\mathbf{Z}_{x}=\sum_{i\in\mathcal{K}}\frac{q_{x}\cdot k_{i}}{\sum_{l\in\mathcal{K}}q_{x}\cdot k_{l}}\cdot v_{i},$ (4)
@@
$\mathbf{Z}_{x}=\sum_{i\in\mathcal{K}}\alpha_{t}(q_{x}\cdot k_{i})\cdot v_{i}=\alpha_{t}q_{x}\cdot\left(\sum_{i\in\mathcal{K}}k_{i}\otimes v_{i}\right).$ (5)
@@
$\mathbf{Z}_{x}=q_{x}+\frac{1}{M+2}\sum_{l=1}^{M+2}\alpha^{l}_{t}q_{x}\cdot\left(\sum_{i_{l}=1}^{N_{l}}k_{i_{l}}\otimes v_{i_{l}}\right),$ (6)
@@
$\mathbf{Z}^{\prime}_{x}=\sum_{i}\alpha_{t}(q_{x}\cdot k_{i})\cdot v_{i}$ (7)
where $q_{x}$, $k_{x}$, and $v_{x}$ are computed from the embedding $\mathbf{Z}_{x}$ as follows:
$q_{x}=W_{q}\mathbf{Z}_{x},\quad k_{x}=W_{k}\mathbf{Z}_{x},\quad v_{x}=W_{v}\mathbf{Z}_{x}.$ (8)